# ConceptWM: Watermarking Diffusion Models for Visual Concept Copyright Protection

### v2

@@
Abstract.
The personalization techniques of diffusion models succeed in generating specific concepts but also pose threats to copyright protection and illegal use. Model Watermarking is an effective method to prevent the unauthorized use of subject-driven or style-driven image generation, safeguarding concept copyrights. However, under the goal of concept-oriented protection, current watermarking schemes typically add watermarks to all images rather than applying them in a refined manner targeted at specific concepts. Additionally, the personalization techniques of diffusion models can easily remove watermarks. Existing watermarking methods struggle to achieve fine-grained watermark embedding with a few images of specific concept and prevent removal of watermarks through personalized fine-tuning. Therefore, we introduce a novel concept-oriented watermarking framework that seamlessly embeds imperceptible watermarks into the concept of diffusion models. We introduce Fidelity-preserving Latent Watermarking (FLW) to generate latent watermarks based on image characteristics and the Adversarial Watermarking Modulation module to prevent ”jailbreaking” via personalized finetuning. To enhance U-Net’s efficiency in learning watermark patterns with limited samples, we propose Efficient Concept Watermark Finetuning, which alternates optimization of model parameters for both watermark embedding and concept learning. We conduct extensive experiments and ablation studies to verify our framework. Our code is available at https://anonymous.4open.science/r/Conceptwm-4EB3/.
Latent Diffusion Model, Image Generation, Model Watermarking
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologies Computer vision
1. Introduction
Recent advancements in diffusion models have significantly upgraded image generation techniques, enabling the creation of realistic images from simple text descriptions and meeting the demand for personalized and customized content. The upgraded diffusion models excel in generating novel concept descriptions, such as unique visual themes, motifs, stylistic preferences, or characters, using only a few images of specific themes or styles (Shi et al., 2024; Gal et al., 2023, 2022; Kumari et al., 2023). Through personalization tools, e.g., DreamBooth (Ruiz et al., 2023) and Lora (Hu et al., 2021), any users with minimal design skills can create images of specific concepts by using only a few personal images and simple text inputs. However, the use of personalization techniques introduces potential risks of concept misuse or leakage. Untrustworthy users may exploit these tools to create unauthorized concepts, leading to violations of original creators’ intellectual property and creativity. For instance, creating concepts related to celebrities or artworks can result in deepfake incidents (Juefei-Xu et al., 2022), the spread of fake news and online spam (Barrett et al., 2023), and serious copyright violations.
@@ 1. Introduction
Figure 1. (a) Post-processing Watermark. (b) U-net Finetuning Watermark. (c) VAE-Finetuning Watermark. (d) Latent-based Watermark. (e) Ours.
Concept watermarking technique is considered to be an alternative to address ownership and accountability issues of specific concepts in generative model, which aims to embed watermark information into the concept and extract the watermarks from subsequent generated images on the basis of the watermarked concept (Feng et al., 2023). In general, the model owner possesses a pre-trained diffusion model and aims to generate images for a new concept. The concept owner provides images of the specific concept to fine-tune the existing diffusion model. During this fine-tuning process, watermark information related to the concept is embedded, enabling the model to generate images with the new concept. A watermark decoder can extract the watermark information from the generated images.
However, as shown in Fig. 1, existing watermarking methods have not yet to achieve concept watermarking, despite the growing interest in adding watermarks to generated images (Kim et al., 2024; Xiong et al., 2023; Zhang et al., 2024; Yang et al., 2024; Fernandez et al., 2023; Zhao et al., 2023b; Wen et al., 2024; Feng et al., 2024; Cox et al., 2007). On the one hand, prior work has rarely addressed fine-grained watermarking mechanisms for specific concepts. Most existing methods (Wen et al., 2024; Fernandez et al., 2023; Zhao et al., 2023b) watermark the entire images rather than images of specific concepts, making them inapplicable for concept watermarking scenario. A major challenge in concept watermarking is to ensure that the watermark not only exists within a specific image but also remains present across various image variations genraeted by the diffusion model. Compared to traditional image watermarks embedded at the pixel level, concept watermarks appears a higher-level technical barrier, as it associates the watermarks with high-level sematic features related to specific concepts. Since the number of images with a specific concept is limited, embedding watermarks in personalization models is a great challenge, especially when fine-tuning the model by injecting a new concept.
On the other hand, robustness remains a challenge for current watermarking methods in concept protection scenarios, when considering diverse adversaries. Recent studies have demonstrated that existing methods are ineffective against model purification attacks (Zhao et al., 2024; Liu et al., 2024), as adversaries can bypass watermarking schemes by generating watermark-free images through personalized fine-tuning. The technical challenge lies in balancing adversariality, image quality, and watermark fidelity. One potential approach to enhancing watermark adversariality is to inject imperceptible protective adversarial perturbations into images, even though this type of methods frequently compromises both image quality and watermark fidelity (Liang et al., 2023; Shan et al., 2023; Van Le et al., 2023). Therefore, there is an urgent need for watermarking solution that ensures concept tracing while effectively addressing adversarial threats.
Our investigation finds three fundamental properties required by model watermarking, including refined nature, robustness, and adversarial properties. To be specific, refined nature, as a core component of diffusion model, requires the ability to embed a watermark into a specific concept through minimal training data and steps without compromising fidelity, as the knowledge of a specific concept generally is learnt by the denoising network (U-Net). Robustness ensures that the watermark remains stable across various image processing operations and generation parameters. Adversarial properties typically guarantee that the generated concept-specific images cannot be easily “jailbreaken” via personalized fine-tuning.
To address these challenges, we propose a novel concept watermarking framework, named ConceptWM, which focuses on two critical factors, namely refined watermark training and adversarial perturbation modulation. To achieve the concept watermark embedding, we develop a Efficient Concept Watermark Finetuning (ECWF) scheme, which rapidly learns watermark embedding patterns in diffusion models with limited data and training resources. Additionally, we develop a Fidelity-preserving Latent Watermark (FLW) module that incorporates the watermark into the latent space and employs a distortion layer in order to enhance robustness against image processing attacks. Next, we propose an Adversarial Watermark Modulation (AWM) to balance adversarial perturbation, watermark fidelity, and image quality, while mitigating the risk of adversarial attacks bypassing the watermark through personalized fine-tuning. This scheme utilizes an alternating optimization strategy to refine both referenced model parameters and adversarial perturbations, so that an optimal equilibrium among adversarial robustness, image quality, and watermark fidelity can be achieved.
@@
Figure 2. Framework of the proposed ConceptWM watermarking scheme for Latent Diffusion Model. The Efficient Concept Watermark Finetuning module employs a two-step training process to learn specific concepts and watermark embedding patterns. The Adversarial Watermark Modulation module utilizes a surrogate model fully fine-tuned on a referenced clean image set $X_{B}$ to guide PGD optimization, iteratively generating adversarial watermarks through multiple rounds of watermark and noise integration.
Our contributions are as follows: (1) We propose a novel concept-driven watermarking framework that seamlessly embeds imperceptible watermarks into the concept of diffusion models. The proposed framework highlights the importance of refined watermark training and fidelity-preserving perturbations in real-world applications of diffusion model watermarking. (2) We propose a mechanism named Adversarial Watermark Modulation that endows concept watermarks with adversarial properties. Through the alternating optimization of model training and adversarial algorithms, it effectively balances adversarial robustness, image quality, and watermark fidelity. (3) Experiments demonstrate that ConceptWM effectively integrates adversarial perturbations into watermarks and maintains better robustness and accuracy in concept watermark embedding and extraction compared with existing 4 baselines over 6.3% image processing attacks.
2. Related Work
@@
Personalization Techniques for Diffusion Models. Recent explorations mainly focus on injecting specific concepts into the model through fine-tunning tools, e.g., Textual Inversion (Gal et al., 2022), DreamBooth (Ruiz et al., 2023) and LoRA (Low-Rank Adaptation)(Hu et al., 2021), given the substaintial computational requirements of training Stable Diffusion from scratch. Specifically, Textual Inversion focuses on training text embedding so that concepts are allowed to be injected into the text encoder without altering weights of the U-Net. Dissimilar to conventioanl text-to-image fine-tuning, DreamBooth fine-tunes the entire U-Net portion of the Stable Diffusion model and employs prior loss during training to prevent overfitting. LoRA is designed to achieve lightweigth efficient fine-tuning by training incremental weights in the attention layers of the U-Net. Our work focuses on watermarking specific concepts generated through DreamBooth fine-tuning.
@@
Model Watermarking for Latent Diffusion Models. The embedding characteristics of different watermarking methods for latent diffusion models are illustrated in Figure 1. It is mainly divided into Post-processing Watermark, U-Net Finetuning Watermark, VAE-Finetuning Watermark, and Latent-based Watermark. These watermarking methods aim to protect the copyright of images or generative models, so they often embed watermarks into all images. For example, WDM (Zhao et al., 2023b) is designed to train an autoencoder to stamp a watermark on all training data before re-training the generator from scratch . Aqualora (Feng et al., 2024) is a lora-based method that watermarks all images. VAE-based methods (Xiong et al., 2023; Kim et al., 2024; Fernandez et al., 2023) fine-tune VAE-Decoder to ensure all generated images containing the watermark. However, this all-image type of methods encounters challenges in protecting specific styles. A large volume of images are generally required for the model to learn the watermark embedding pattern, but concept-oriented images for finetuning are limited to only a few.
@@
Protective Perturbation against Stable Diffusion. Recent studies mostly have addressed disrupting the fine-tuning process by adding imperceptible protective noise to images, in order to prevent personal images from potential infringement during the fine-tuning of Stable Diffusion. For example, AdvDM employs adversarial noise to safeguard personal images on Stable Diffusion by maximizing the Mean Squared Error loss during optimization (Liang et al., 2023). Anti-DreamBooth (Van Le et al., 2023) offers a bi-level min-max optimization for generating protective perturbations, through integrating the DreamBooth fine-tuning process of Stable Diffusion into its framework. Other studies (Wu et al., 2023; Ye et al., 2023; Zhao et al., 2023a; Zheng et al., 2023) have explored similar adversarial perturbation techniques to generate protective noise for images.
3. Methodology
3.1. Overview
@@
The main objective of our scheme is to protect images of a specific concept that are generated by the diffusion model, by embedding an adversarial watermark pattern into the U-Net component of the diffusion model. Fig. 2 illustrates an overview of our approach. To embed a watermark that is both fidelity-preserving and conducive to learning by U-Net, we develop a Fidelity-preserving Latent Watermarking (FLW) module, which adds a combination layer for the message encoder and latent concatenation, so that watermarks can be generated based on image characteristics rather than a fixed watermark signal. To protect the generated concept-specific images from being “jailbroken” through personalized finetuning, we develop an Adversarial Watermarking Modulation module to integrate adversarial propeties into the watermark. To enable U-Net to efficiently learn watermark patterns during concept learning with limited samples, we present an Efficient Concept Watermark Finetuning module that alternates optimization of model parameters to simultaneously achieve adversarial watermark embedding and concept learning.
3.2. Fidelity-preserving Latent Watermark
Figure 3. The comparison of two variants of Latent watermark namely Overlay-based Latent Watermarking and Fusion-based Latent Watermarking.
@@ 3.2. Fidelity-preserving Latent Watermark
Existing image watermarking techniques frequently arise substantial discrepancies between the watermark embedding pattern and the latent space. This mismatch issue may cause the disruption or even complete loss of watermark information when the image is encoded into the latent space by a VAE encoder or regenerated through diffusion, as has been discussed by recent watermarking attacks (An et al., 2024; Ballé et al., 2018; Cheng et al., 2020). Thus, we aim to embed a robust watermark within the image that is both resilient and conducive to efficient learning for U-Net. To be specific, we have developed two variants of latent watermarking, which are Overlay-based Latent Watermarking (OLW) and Fidelity-preserving Latent Watermarking (FLW). As shown in Fig. 3, both variants involve the latent domain perturbations. The watermarked images are processed through the message decoder to evaluate the discrepancy between the decoded watermark information and the original content of the watermarked image. The message loss is the Binary Cross Entropy (BCE) between the message $m$ and $\sigma(m^{\prime})$. Given the message length $N$, the message loss $\mathcal{L}_{m}$ is defined in Eq. (1).
(1) $\mathcal{L}_{m}=-\sum_{k=1}^{N}m_{k}\log\sigma(m_{k}^{\prime})+(1-m_{k})\log(1-\sigma(m_{k}^{\prime}))$
@@
Fidelity-preserving Latent Watermarking. To address issues of OLW, we have developed Fidelity-preserving Latent Watermarking (FLW) that adds a combination layer after the message encoder and latent concatenation. To ensure visual consistency, we calculate LPIPS loss $\mathcal{L}_{pips}$ (Zhang et al., 2018) and utilize a Peak Pixel Difference (PPD) Loss $\mathcal{L}_{ppd}$ which selects the most prominent parts of the image and calculates channel loss to reduce excessive modification of the images. Our training objective ($\mathcal{L}_{total}$) can be expressed in Eq. (2), where $\lambda$ and $\mu$ represent the coefficients.
(2) $\mathcal{L}_{total}=\mathcal{L}_{m}+\lambda\mathcal{L}_{pips}+\mu\mathcal{L}_{ppd}$
The jointly optimized message encoder and decoder, working in concert with the Adversarial Watermark module, will empower the U-Net to learn robust watermarking patterns.
3.3. Adversarial Watermark Modulation
To protect the generated concept-specific images from “jailbreaking” through personalized fine-tuning, we integrate adversarial properties into the watermark. By maximizing the image generation loss with adversarial noise added to the watermark and minimizing the concept learning loss through Dreambooth fine-tuning with reference images, we ensure that the adversarial watermark does not interfere with concept learning while preventing “jailbreaking” through personalized fine-tuning.
@@ 3.3. Adversarial Watermark Modulation
Algorithm 1 Adversarial Watermark Modulation
1:experiment AWM(($x_{0}$,$X_{B}$,$\theta$))
2:     for  $x_{b}\in X_{B}$ do
3:         c= “ a photo of $\langle$ concept $\rangle$ person ”
4:         $\theta=\operatorname{TrainOneStepDreamBooth}(\theta,x_{b},c)$
5:     end for
6:     $s\leftarrow secret$
7:     $x_{w}=LatentWatermarkEmbedding(x_{0},s)$
8:     $\sigma_{1}=x_{w}-x_{0}$
9:     $Initialize$ $x_{0}^{0}\leftarrow x_{0}$
10:     for $t=1\ldots T$ do
11:         $\sigma_{2}^{*(i)}=\alpha\cdot sign(\nabla_{x}\mathcal{L}_{cond}(\theta,x_{0}^{t}+\sigma_{1}))$
12:         ${x_{0}}^{t+1}={x_{0}}^{t}+\sigma_{2}^{*(i)}$
13:     end for
14:     $\varphi=x_{0}^{T}+\sigma_{1}-x_{0}$
15:return $\varphi$
Adversarial Watermark Modulation. The goal of this modulation is to maximize $L_{cond}$ in Eq. (3) to hinder personalized finetuning and minimize $L_{ft}$ in Eq. (4) to learn the concept. Our solution is to use watermark-free images $\mathcal{X}_{A}$ for one-step DreamBooth training to reduce the DreamBooth loss $\mathcal{L}_{\text{ft}}$, while simultaneously applying Projected Gradient Descent (PGD) (Madry, 2017) to enhance the adversarial loss $L_{cond}$. The model used for PGD is trained with clean referenced images $\mathcal{X}_{B}$ in a one-step dreambooth training process, rather than the images used for concept learning. The purpose of this approach is to allow the model to learn adversarial perturbations without affecting the learning of the concept. The specific procedure is defined by Eq.s (5) and (6), where $\sigma_{1}$ denotes the residual of the watermarked image and original image. $x^{i}$ denotes the clean samples.
(5) $\theta^{*}=\arg\min_{\theta}\sum_{x\in\mathcal{X}_{A}}\mathcal{L}_{ft}(\theta,x^{i})$
(6) $\varphi^{(i)}=\arg\max_{\sigma_{2}^{(i)}}\mathcal{L}_{cond}(\theta^{*},x^{(i)}+\sigma_{1}+\sigma_{2}^{(i)})$
The adversarial noise $\sigma^{*(i)}$ is within the constraint range of $\eta$. The Adversarial Watermark Modulation outputs an adversarial watermark $\varphi$ as the parameter for Concept-oriented Watermark Fine-tuning to facilitate learning.
Algorithm 2 Efficient Concept Watermark Finetuning
1:experiment $CWF$($X_{A}$,$X_{B}$,$\theta$,E)
2:     for  $x_{0}\in X_{A}$ do
3:         c=“ a photo of $\langle$ concept $\rangle$ person ”
4:         $\theta=\operatorname{TrainOneStepDreamBooth}(\theta,x_{0},c)$
5:         $\theta^{*}\leftarrow\text{ copy parameters from }\theta$
6:         $\varphi=AWM(x_{0},X_{B},\theta)$
7:         $t\sim\mathrm{Uniform}(\{1,\ldots,T\})$
8:         $z_{0}=E(x_{0})$
9:         $z_{t}\leftarrow\sqrt{\bar{\alpha}_{t}}z_{0}+\sqrt{1-\bar{\alpha}_{t}}\boldsymbol{\epsilon}$
10:         $z_{t}^{\mathrm{w}}\leftarrow\sqrt{\bar{\alpha}_{t}}(z_{0}+\varphi)+\sqrt{1-\bar{\alpha}_{t}}\boldsymbol{\epsilon}$
11:         Take gradient step on $\nabla_{\theta}\|{\epsilon_{\theta}}(z_{t}^{\mathrm{w}},t,c)-{\epsilon_{\theta^{*}}}(z_{t},t,c)\|^{2}$
12:     end for
13:return $\theta$
@@
3.4. Efficient Concept Watermark Finetuning
The training objective of this stage is defined by Eq. (7), where $\varphi$ represents the adversarial perturbation containing the watermark generated through AWM, $\epsilon_{\theta}^{*}$ denotes the fixed model, and $\epsilon_{\theta}$ refers to the trained model.
(7) $\mathcal{L}_{\mathrm{cwft}}(\theta,x_{t})$ $=\mathbb{E}_{t,c,x}\Big{[}\|\boldsymbol{\epsilon}_{\theta}(x_{t}+\varphi,t,c)-\boldsymbol{\epsilon}_{\theta}^{*}(x_{t},t,c\big{)}\|^{2}\Big{]}$
Algorithm 2 outlines the finetuning procedure for the U-Net denoising model $\theta$ with three components: target concept images $X_{A}$, clean reference images $X_{B}$, and the VAE encoder $E$. The process has three phases: In concept specialization phase (Lines 1-5), The model $\theta$ is fine-tuned on $X_{A}$ to produce a specialized model $\theta^{*}$ that captures the target concept while maintaining general denoising capabilities. In adversarial watermark generation phase (Line 6), an adversarial perturbation is computed through the Fidelity-preserving Latent Watermark in section 3.2. In watermark finetuning phase (Lines 7-13), the model is fed with both watermarked and non-watermarked images, and the difference in the predictions is used as a loss to guide the generative model in learning the watermark embedding pattern.
[…]
Figure 4. Demonstration of the generated images. Below each image, we display the corresponding Bit Accuracy, along with the image quality metrics PIQE and NIQE.
4. Experiment
4.1. Implementation Details
Datasets. To evaluate the effectiveness of the proposed methods, for subject-driven generation, we select the face dataset CelebA-HQ (Zhu et al., 2022). For style-driven generation, we select 12 artists with distinct styles from WikiArt  [^1] https://www.wikiart.org, with each artist contributing 10 artworks that are consistent in terms of both artistic style and time period. Each image is resized and center-cropped to a resolution of 512 × 512. In each experiment, a total of 5–8 carefully selected images were utilized for DreamBooth fine-tuning.
[…]
Type Method Concept Oriented Concept Protection Fidelity Robustness
FID $\uparrow$ CLIP $\uparrow$ BitAcc $\uparrow$ BitAcc(Adv.) $\uparrow$ TPR $\uparrow$ TPR (Adv.$\uparrow$
@@ 4.1. Implementation Details
Post-Processing DwtDctSvd $\times$ $\times$ 25.01 0.363 100.0 71.47 1.000 0.367

### v3

# Watermarking Visual Concepts for Diffusion Models

@@
Abstract
The personalization techniques of diffusion models succeed in generating images with specific concepts. This ability also poses great threats to copyright protection and network security since malicious users can generate unauthorized content and disinformation relevant to a target concept. Model watermarking is an effective solution to trace the malicious generated images and safeguard their copyright. However, existing model watermarking techniques merely achieve image-level tracing without concept traceability. When tracing infringing or harmful concepts, current approaches execute image concept detection and model tracing sequentially, where performance is critically constrained by concept detection accuracy. In this paper, we propose a lightweight concept watermarking framework that efficiently binds target concepts to model watermarks, supporting simultaneous concept identification and model tracing via single-stage watermark verification. To further enhance the robustness of concept watermarking, we propose an adversarial perturbation injection method collaboratively embedded with watermarks during image generation, avoiding watermark removal by model purification attacks. Experimental results demonstrate that ConceptWM significantly outperforms state-of-the-art watermarking methods, improving detection accuracy by 6.3%-19.3% across diverse datasets including COCO and StableDiffusionDB. Additionally, ConceptWM possesses a critical capability absent in other watermarking methods: it sustains a 21.7% FID/CLIP degradation under adversarial fine-tuning of Stable Diffusion models on WikiArt and CelebA-HQ, demonstrating its capability to mitigate model misuse.
Codehttps://anonymous.4open.science/r/Conceptwm
Introduction
Recent advance in diffusion models have significantly improved image generation ability, enabling the creation of high-quality images from personalized textual descriptions. The upgraded diffusion models achieve few-shot learning for new visual concept generation, such as unique visual themes, motifs, stylistic preferences, or celibrities, using only a few images with the target concepts (Shi et al. 2024; Gal et al. 2023, 2022; Kumari et al. 2023). However, these techniques pose significant risks for visual concept infringement and misuse. For example, a malevolent actor might leverage a few online trademark or digital artwork images to fine-tune off-the-shelf diffusion models and edit the content without permission (Zhu, Takahashi, and Kataoka 2024; Li et al. 2025). Similarly, a few online celebrity images are adequate for malevolent actors to generate misinformation with these celebrities (Wang et al. 2024). It becomes an urgent issue to effectively trace AI-edited visual concept images for copyright protection and privacy preservation.
@@ Introduction
Figure 1: Five main kinds of diffusion model watermarking methods: (a) Post-processing watermark. (b) VAE finetuning watermark. (c) U-Net finetuning watermark. (d) Latent-based watermark. (e) Concept-oriented watermark.
Model watermarking is a representative way to trace the source model of generated images and thus verify the accountability of malicious content. The typical solutions of model watermarking fine-tune the pre-trained diffusion models to inject watermarks and extrect the watermark information from any generated images by a watermark decoder. Compared with post processing solutions by adding watermarks after image generation (Fig. 1(a)), model watermarking is more secure to avoid watermark evasion (Zhao et al. 2023b). Despite the great progress in diverse model watermarking schemes (Fig. 1(b)(c)(d)), existing schemes indiscriminately inject watermarks into all the generated images agnostic to visual concepts, which has two obvious limitations for visual concept tracing and protection. On the one hand, tracing concept-level malicious image becomes indirect that requires a two-stage pipeline: initial detection of target visual concepts in images, followed by watermark-based tracing for concept images. Thus, the tracing effectiveness is critically constrained by concept detection accuracy. One the other hand, recent studies (Cui et al. 2025) reveal that existing personalization methods enable malicious actors to remove watermarks by fine-tuning diffusion models with a few watermarked samples, while retaining concept generation capabilities. Therefore, current model watermarking methods with only image tracing capability are still unable to prevent misue of watermarked images.
In this paper, we propose an approach that aims to trace and protect generated images with target visual concepts by diffusion model. The core idea of our approach is Concept Watermarking, that is, during image generation, we embed both traceable watermarks and fine-tuning-resistent perturbations into concept-specific generated images. This paradigm enables people to technically govern AI-powered malicous content and proactively defend subsequent misuse of generated content in a finer manner. To this end, three new challenges need to be addressed: First, for a target concept (e.g., ‘a specific kind A of cats’), concept watermarking requires to establish tripartite correlation among a concept watermark (e.g., a bit sequence ‘01100’), visual concepts (diverse appreances of ‘A cats’) in the generated images, and a textual concept (a psuedo word of ‘cat A’) in the prompts, which is quite different from existing watermarking methods that solely establish bipartite correlation between a watermark and model outputs; Second, concept watermarking has adversarial capability that the generated concept images are unable to be used for model fine-tuning, which has no influence on concept-irrelevant generated images; Third, concept watermarking is robust to diverse real-world attacks including image processing, geometric transformations, and model regeneration attacks.
To address these challenges, we propose a lightweight concept watermarking framework, named ConceptWM, which contains two main componets, i.e., efficient concept watermark training and imperceptible adversarial perturbation injection. To embed a watermark that is both fidelity-preserving and conducive to learning by U-Net, we propose a fidelity-preserving latent watermark module that incorporates the watermark into the latent space and introduce a fusion layer after the message encoder and latent concatenation to mitigate rigid watermark pattern artifacts. To achieve concept-level image tracing, we propose an efficient concept watermark training strategy that alternately optimizes DreamBooth fine-tuning and watermark embedding training to achieve joint concept-watermark learning. Our method learns watermark embeddings efficiently, requiring only a few reference images and low training budgets. To prevent the misuse of watermarked images, we propose an imperceptible adversarial perturbation injection module to refine both referenced model parameters and adversarial perturbations by proposed optimization strategy, which balances adversarial robustness, watermark fidelity, and image quality, while mitigating the attack risk by personalized fine-tuning.
Our contributions are as follows: (1) We propose a lightweight concept watermarking framework that imperceptibly embeds watermarks into the generated concept-specific images by diffusion models. Concept watermarking first tackles the challenge of establishing tripartite correlation across three modalities, e.g., watermarks, visual concept and textual concept, which makes it possible for concept-level generated image tracing in one-stop manner. (2) We propose a new adversarial mechanism that endows concept watermarks with proactive denfence capabilities. This mechanism enables the adversarial perturbation injection without interference to image quality and watermark fidelity, while making the generated concept images resistent to malicious use for model fine-tuning. (3) ConceptWM achieves 6.3% higher robustness against image processing than 4 baselines while maintaining a critical property: it forces 21.7% FID/CLIP degradation in fine-tuned Stable Diffusion models - a capability beyond current methods.
@@
Figure 2: Framework of ConceptWM. The ECWT module employs a two-step training process to learn specific concepts and watermark patterns. The IAPI module utilizes a surrogate model fully fine-tuned on a referenced clean image to guide PGD optimization, iteratively generating adversarial watermarks through multiple rounds of watermark and noise integration.
Related Work
@@
Personalization Techniques for Diffusion Models. Recent explorations mainly focus on injecting specific concepts into the model through fine-tunning tools, e.g., Textual Inversion (Gal et al. 2022), DreamBooth (Ruiz et al. 2023) and LoRA (Low-Rank Adaptation) (Hu et al. 2021). Specifically, Textual Inversion focuses on training text embedding so that concepts are allowed to be injected into the text encoder without altering weights of the U-Net. DreamBooth fine-tunes the entire U-Net portion of the Stable Diffusion model and employs prior loss during training to prevent overfitting. LoRA is designed to achieve lightweigth efficient fine-tuning by training incremental weights in the attention layers of the U-Net. Our work focuses on watermarking specific concepts generated through DreamBooth fine-tuning.
@@
Model Watermarking for Latent Diffusion Models. The embedding characteristics of different watermarking methods for diffusion models are illustrated in Figure 1, These watermarking methods aim to protect the copyright of images or generative models, so they often embed watermarks into all images. WDM (Zhao et al. 2023c) is designed to train an autoencoder to stamp a watermark on all training data before re-training the generator from scratch. Aqualora (Feng et al. 2024) is a lora-based method that watermarks all images. VAE-based methods (Xiong et al. 2023; Kim et al. 2024; Fernandez et al. 2023) fine-tune VAE-Decoder to ensure all generated images containing the watermark. However, these methods fail to protect specific concepts. Our Efficient Concept Watermark Training embeds watermark into specific concept with a few images of specific concept.
@@
Protective Perturbation against Stable Diffusion. Recent studies (Wu et al. 2023; Ye et al. 2023; Zhao et al. 2023d; Zheng et al. 2023) mostly have addressed disrupting the fine-tuning process by adding imperceptible protective noise to images, in order to prevent personal images from potential infringement during the fine-tuning of Stable Diffusion. AdvDM employs adversarial noise to safeguard personal images on Stable Diffusion by maximizing the Mean Squared Error loss during optimization (Liang et al. 2023). Anti-DreamBooth (Van Le et al. 2023) offers a bi-level min-max optimization for generating protective perturbations, through integrating the DreamBooth fine-tuning process of Stable Diffusion into its framework. Our Imperceptible Adversarial Perturbation Injection effectively endows the watermark with the aforementioned adversarial properties.
Methodology
@@
The main objective of our scheme is to protect images of a specific concept that are generated by the diffusion model, by embedding an adversarial watermark pattern into the U-Net component of the diffusion model. Fig. 2 illustrates an overview of our approach. To embed a watermark that is both fidelity-preserving and conducive to learning by U-Net, we develop a Fidelity-preserving Latent Watermarking (FLW) module, which adds a combination layer for the message encoder and latent concatenation, so that watermarks can be generated based on image characteristics rather than a fixed watermark signal. To protect the watermarked concept-specific images from misusing, we develop an Imperceptible Adversarial Perturbation Injection module to integrate adversarial propeties into the watermark. To enable U-Net to efficiently learn watermark patterns during concept learning with limited samples, we present an Efficient Concept Watermark Finetuning module that alternates optimization of model parameters to simultaneously achieve adversarial watermark embedding and concept learning.
Fidelity-preserving Latent Watermark
@@
Existing image watermarking techniques frequently arise substantial discrepancies between the watermark embedding pattern and the latent space. This mismatch issue may cause the disruption or even complete loss of watermark information when the image is encoded into the latent space by a VAE encoder or regenerated through diffusion, as has been discussed by recent watermarking attacks (An et al. 2024; Ballé et al. 2018; Cheng et al. 2020). Thus, we aim to embed a robust watermark within the image that is both resilient and conducive to efficient learning for U-Net. To be specific, we have developed two variants of latent watermarking, which are Overlay-based Latent Watermarking (OLW) and Fidelity-preserving Latent Watermarking (FLW). Both variants involve the latent domain perturbations. The watermarked images are processed through the message decoder to evaluate the discrepancy between the decoded watermark information and the original content of the watermarked image. The message loss is the Binary Cross Entropy (BCE) between the message $m$ and $\sigma(m^{\prime})$. Given the message length N, the message loss $\mathcal{L}_{m}$ is defined in Eq. (1).
$\mathcal{L}_{m}=-\sum_{k=1}^{N}m_{k}\log\sigma(m_{k}^{\prime})+(1-m_{k})\log(1-\sigma(m_{k}^{\prime}))$ (1)
@@
Fidelity-preserving Latent Watermarking. To address issues of OLW, we have developed Fidelity-preserving Latent Watermarking (FLW) that adds a combination layer after the message encoder and latent concatenation. To ensure visual consistency, we calculate LPIPS loss $\mathcal{L}_{pips}$ (Zhang et al. 2018) and utilize a Peak Pixel Difference (PPD) Loss $\mathcal{L}_{ppd}$ which selects the most prominent parts of the image and calculates channel loss to reduce excessive modification of the images. Our training objective ($\mathcal{L}_{total}$) can be expressed in Eq. (2), where $\lambda$ and $\mu$ represent the coefficients.
$\mathcal{L}_{total}=\mathcal{L}_{m}+\lambda\mathcal{L}_{pips}+\mu\mathcal{L}_{ppd}$ (2)
The jointly optimized message encoder and decoder, working in concert with the Adcersarial Perturbation Modulation module, will empower the U-Net to learn robust watermarking patterns. We provide the training structure comparison of OLW and FLW in the Appendix.
Imperceptible Adversarial Perturbation Injection
To protect the generated concept-specific images from removing through personalized fine-tuning, we integrate adversarial properties into the watermark. By maximizing the image generation loss with adversarial noise added to the watermark and minimizing the concept learning loss through Dreambooth fine-tuning with reference images, we ensure that the watermark does not interfere with concept learning while preventing the misuse of watermarked image.
@@
Adversarial Perturbation Injection. The goal of this modulation is to maximize $L_{cond}$ in Eq. (3) to hinder personalized finetuning and minimize $L_{ft}$ in Eq. (4) to learn the concept. Our solution is to use watermark-free images $\mathcal{X}_{A}$ for one-step DreamBooth training to reduce the DreamBooth loss $\mathcal{L}_{\text{ft}}$, while simultaneously applying Projected Gradient Descent (PGD) (Madry 2017) to enhance the adversarial loss $L_{cond}$. The model used for PGD is trained with clean referenced images $\mathcal{X}_{B}$ in a one-step dreambooth training process, rather than the images used for concept learning. The purpose of this approach is to allow the model to learn adversarial perturbations without affecting the learning of the concept. The specific procedure is defined by Eq.s (5) and (6), where $\sigma_{1}$ denotes the residual of the watermarked image and original image. $x^{i}$ denotes the clean samples.
$\theta^{*}=\arg\min_{\theta}\sum_{x\in\mathcal{X}_{A}}\mathcal{L}_{ft}(\theta,x^{i})$ (5)
$\varphi^{(i)}=\arg\max_{\sigma_{2}^{(i)}}\mathcal{L}_{cond}(\theta^{*},x^{(i)}+\sigma_{1}+\sigma_{2}^{(i)})$ (6)
The adversarial noise $\sigma^{*(i)}$ is within the range of $\eta$. The Imperceptible Adversarial Perturbation Modulation outputs an adversarial watermark $\varphi$ as the parameter for Concept-oriented Watermark Training. The pseudo-code of Adversarial Perturbation Modulation is presented in the Appendix.
Efficient Concept Watermark Training
The training objective of this stage is defined by Eq. (7), where $\varphi$ represents the adversarial perturbation containing the watermark generated through IAPI, $\epsilon_{\theta}^{*}$ denotes the fixed model, and $\epsilon_{\theta}$ refers to the trained model.
$\mathcal{L}_{\mathrm{cwft}}(\theta,x_{t})$ $=\mathbb{E}_{t,c,x}\Big{[}\|\boldsymbol{\epsilon}_{\theta}(x_{t}+\varphi,t,c)-\boldsymbol{\epsilon}_{\theta}^{*}(x_{t},t,c\big{)}\|^{2}\Big{]}$ (7)
The pseudo-code of Concept-oriented Watermarking Training is presented in the Appendix.
@@
Figure 3: Demonstration of the generated images. Below each image, we display the corresponding Bit Accuracy, along with the image quality metrics PIQE and NIQE.
Type Method
Concept
Tracing
Concept
Protection
Fidelity Robustness
FID CLIP BitAcc BitAcc(Adv.) TPR TPR (Adv.)
@@
Post-Processing DwtDctSvd $\times$ $\times$ 25.01 0.363 99.98 71.47 1.000 0.367